OpenStudy (anmsteelers):

Explain the distance formula. Then use it for calculate the distance between A(1,1) and B(7,-7). Can someone help?

OpenStudy (will.h):

do you know the distance formula?

OpenStudy (anmsteelers):

no

OpenStudy (will.h):

\[\sqrt{(y2-y1)^2 + (x2-x1)^2}\] Apply the data you have and you'll get your answer.

OpenStudy (will.h):

need any additional help?

OpenStudy (anmsteelers):

yes

OpenStudy (will.h):

okay. \[\sqrt{(y2-y1)^2 + (x2-x1)^2}\] We have the coordinates (1,1) (7,-7) so just really plug them in the formula above. \[\sqrt{(-7 - 1)^2 + (7 - 1)^2}\] So the answer would be 10 Hope that helps

OpenStudy (anmsteelers):

Thanks!